Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x.

f(x) = x2 - 3 and g(x) = square root of quantity three plus x
1
Expert's answer
2020-09-23T16:50:46-0400

"f(x) = x^2 - 3, g(x) = \sqrt{3 + x}\\


f(g(x)) \hspace{0.1cm} \textsf{means}\hspace{0.1cm}\textit{ f of g},\\
\textsf{i.e insert the function}\hspace{0.1cm} g\\\textsf{which is expressed as a function of }\hspace{0.1cm} x\\ \textsf{in the}\hspace{0.1cm} x \hspace{0.1cm}\textsf{in function} \hspace{0.1cm}f. \\

f(g(x)) = (\sqrt{3 + x})^2 - 3\\


f(g(x)) = 3 + x - 3 = x\\



\textsf{similarly}\hspace{0.1cm}g(f(x)) = \sqrt{3 + (x^2 - 3)} = \sqrt{3 + x^2 - 3} = \sqrt{x^2} = x\\


\therefore \textsf{Since}\hspace{0.1cm} f(g(x)) = g(f(x)) = x,\\
\textsf{functions} \hspace{0.1cm}f \hspace{0.1cm}\textsf{and} \hspace{0.1cm}g\hspace{0.1cm}\textsf{are inverses of one another.}"